OpenStudy (cutiecomittee123):

log*bA=3 log*bC=2 log*bD=5 log*b A^3D^4/C^2

OpenStudy (cutiecomittee123):

Oh yeah what i meant to add was that we are trying to figure out the value of the second eqution given the values of the a c and d

OpenStudy (cutiecomittee123):

also the options for answers are 176 25 33 and 4218.75

OpenStudy (campbell_st):

well apply the log laws for multiplication and addition log law for division is subtract \[\log_{b} (\frac{A^3 D^4}{C^2}) = \log_{b} (A^3D^4) - \log_{b} (C^2)\] then you need the log law for multiplication, which is add the logs \[\log(AB) = \log(A) + \log(B)\] so apply this law to the A^3D^4 lastly you need to apply the log law for powers... which needs you to multiply the log by the power \[\log(x^a) = a \times \log(x)\] if you can do all that, the last thing will be to substitute the values that have been given and calculate an answer.

OpenStudy (cutiecomittee123):

So confused! pls explain more:)

OpenStudy (cutiecomittee123):

@campbell_st

OpenStudy (campbell_st):

ok... so the next step is \[\log_{b}(A^3) + \log_{b}(D^4) - \log_{b}(C^2)\] does that make sense...?

OpenStudy (cutiecomittee123):

Yes I did that:)

OpenStudy (cutiecomittee123):

@campbell_st now what

OpenStudy (cutiecomittee123):

@kenzie.christiansen do you know how to do this

OpenStudy (kenzie.christiansen):

no sorry if i did i could help u but i dont

OpenStudy (campbell_st):

ok...apply the log law for powers \[3 \times \log_{b}(A) + 4 \times \log_{b}(D) - 2 \times \log_{b}(C)\] now substitute your values.